# Not Just How Much, But Where: Decomposing Epistemic Uncertainty into Per-Class Contributions

### v1

@@ 2 Per-Class Epistemic Uncertainty
We decompose MI into a per-class vector via a second-order Taylor expansion of the entropy function (Sections 2.12.2), show that the decomposition corrects a pathology of raw variance near the simplex boundary (Section 2.3), characterise its axiomatic properties (Section 2.4), and introduce a diagnostic that flags when the approximation degrades (Section 2.5).
@@
with diagonal elements $\mathrm{Var}[p_{k}]=\frac{1}{S-1}\sum_{s}(p_{k}^{(s)}-\mu_{k})^{2}$. Throughout we rely on the standard decomposition
@@ 2.1 From Scalar MI to a Per-Class Decomposition
The Hessian is diagonal and negative definite on the interior of the simplex. The diagonal Hessian is what makes a per-class decomposition possible: there are no cross-class terms at second order, so the quadratic approximation splits cleanly over classes.
Theorem 2.2 (MI Approximation).
@@
The simplex constraint forces $\mathrm{Cov}[\boldsymbol{p}]$ to rank K1K{-}1, coupling the $K$ components; this does not invalidate the trace formula. See Appendix B.1 for proof. 222smith2018understanding also exploit a Taylor expansion to relate softmax variance to MI, but expand the logarithm $\log p_{k}$ rather than the entropy function $H(\boldsymbol{p})$ itself (their Section 3.2, Eq. 10). This yields a class-averaged variance $\frac{1}{K}\sum_{k}\mathrm{Var}[p_{k}]$ with equal weights per class, whereas expanding the entropy via its Hessian yields the curvature-weighted sum $\frac{1}{2}\sum_{k}\mathrm{Var}[p_{k}]/\mu_{k}$, in which the $1/\mu_{k}$ factor enables per-class attribution..
@@ 2.3 Why Variance Alone Fails: Boundary Suppression
which approaches $\frac{1}{2}$ (not zero) as $\mu_{k}\to 0$. The contrast is stark: while $\mathrm{Var}[p_{k}]$ is crushed to zero near the simplex boundary, $C_{k}$ retains a non-vanishing upper bound. See Appendix B.2.
@@ 3.1 Experimental Setup
Uncertainty is estimated from $S=30$ stochastic forward passes through the posterior. Our primary safety metric is the critical false negative rate (Critical FNR): among critical-class samples retained at a given coverage level, the fraction misclassified as safe. We summarise performance across coverage levels via the area under the selective risk curve (AUSC) [elyaniw2010], with 95% CIs from 200 bootstrap resamples. On the test set, the Bayesian model achieves accuracy $0.8$ and quadratic weighted kappa $0.65$; per-class performance in Appendix D.3.
@@ 3.3 Results
Table 2 reports AUSC and Critical FNR at 80% coverage with bootstrap CIs. $C_{\text{crit\_max}}$ achieves the lowest AUSC ($0.285$), reducing the area under the selective risk curve by 34.7% relative to MI ($0.436$, $p<0.005$) and by 56.2% relative to Sale_EU_crit ($0.650$, $p<0.005$). At the clinically relevant 80% coverage operating point, $C_{\text{crit\_max}}$ achieves FNR $0.302$, compared to $0.339$ for MI and $0.409$ for Sale_EU_crit.
@@
Figure 1 visualises these results: $C_{\text{crit\_max}}$ dominates across the full coverage range (left), with non-overlapping bootstrap interquartile ranges against scalar baselines (right).
@@
The improvement over raw variance is especially striking. Sale_EU_crit and Var_crit (AUSC $0.650$ and $0.606$) perform worse than entropy ($0.604$), confirming that boundary suppression (Lemma 2.4) severely confounds unnormalised variance for critical classes with $\mu_{k}\approx 0.06$. The $1/\mu_{k}$ normalisation in $C_{k}$ corrects precisely this: Grade 3 has mean probability $\sim$0.06, giving entropy curvature $\sim$17, so even small variance there carries large information-theoretic weight. This normalisation benefit is robust to the choice of inference method: a deep ensemble [lakshminarayanan2017simple] of five members yields $C_{\text{crit\_max}}$ AUSC $0.390$ vs. $0.447$ for Sale_EU_crit ($12.9\%$ reduction; Appendix D.12).
@@
MC dropout comparison.
To test robustness to the posterior approximation, we repeat the evaluation using standard MC dropout [gal2016dropout] (p=0.3p{=}0.3, S=30S{=}30). The ranking changes markedly: CBEC achieves the lowest AUSC ($0.197\pm 0.012$, winning 100% of bootstrap samples), a 53.6% reduction over MI ($0.425\pm 0.024$), while $C_{\text{crit\_max}}$ drops to $0.419$. MC dropout is known to produce regions of spurious confidence where the approximate posterior underestimates uncertainty [smith2018understanding]. In our framework, this inflates $\rho_{k}$ for critical classes, degrading the Taylor approximation underlying $C_{k}$ while leaving CBEC’s correlation gate intact. This confirms the intended complementarity: $C_{\text{crit\_max}}$ is the primary metric under well-calibrated posteriors; CBEC provides a robust fallback when $\rho_{k}$ signals degradation. Full results in Appendix D.11.
@@ 4.1 Setup
We consider two OoD detection tasks spanning different modalities and class structures.
FashionMNIST [xiao2017fashion] $\to$ KMNIST [clanuwat2018deep]: A 10-class image classification task where the model is trained on FashionMNIST (50,000 train, 10,000 test) and evaluated on KMNIST (10,000 samples) as OoD.
MIMIC-III ICU [johnson2016mimic] $\to$ Newborn: A binary mortality prediction task on tabular clinical data (44 features) where the model is trained on adult ICU admissions (40,406 train, 4,490 test) and evaluated on the Newborn unit (5,357 samples) as OoD.
@@ 4.2 Results
Table 3 summarises OoD detection performance. $\sum_{k}C_{k}$ achieves the highest AUROC on both datasets: $0.735\pm 0.009$ on FashionMNIST versus $0.724$ for MI, and $0.815\pm 0.017$ on MIMIC-III versus $0.802$ for MI. The OoD-to-ID ratio column reveals the mechanism: on FashionMNIST, $\sum_{k}C_{k}$ amplifies the distributional shift signal to $6.43\times$ versus $5.92\times$ for MI. On FashionMNIST, this effect is predicted by Lemma 2.9: the gap $\sum_{k}C_{k}-\text{MI}\approx\frac{1}{6}\sum_{k}m_{3,k}/\mu_{k}^{2}$ grows with posterior skewness, and OoD samples produce higher skewness diagnostic $\rho_{k}$ than ID samples across all ten classes (Appendix E.6, Figure 14). On MIMIC-III the picture is more nuanced: the survival class (k=0k{=}0) follows the expected pattern ($\rho_{0}^{\text{OOD}}>\rho_{0}^{\text{ID}}$, median shift $43\%$), but the mortality class (k=1k{=}1) reverses because near-zero $\mu_{1}$ for newborns collapses $\mathrm{Var}[p_{1}]$ faster than $1/\mu_{1}$ can compensate, concentrating the third-order signal entirely in class 0 (Appendix E.6).
The unnormalised baseline $EU_{\text{var}}$ underperforms MI on both datasets despite matching its ratio on FashionMNIST ($5.92\times$): without mean-normalisation, raw variance $\sigma^{2}_{k}$ is bounded above by $\mu_{k}(1-\mu_{k})$, compressing the dynamic range of uncertainty scores. Strikingly, on MIMIC-III $EU_{\text{var}}$ achieves the highest ratio ($1.71\times$) yet the lowest AUROC among epistemic measures. A high mean ratio does not imply good separability: AUROC depends on the full distributional overlap, and the un-normalised variance sum concentrates both ID and OoD scores near zero with a shared heavy tail, so the numerically large mean shift is swamped by within-group spread (Appendix E.5, Figure 12). This pattern is consistent across inference methods: a deep ensemble replicates the same ranking ($\sum_{k}C_{k}{=}0.753>\text{MI}{=}0.752>EU_{\text{var}}{=}0.750$) on MIMIC-III, with $EU_{\text{var}}$ again achieving the highest ratio ($2.72\times$) and the lowest AUROC (Appendix E.7).
@@
On the binary MIMIC-III task, the decomposition $\sum_{k}C_{k}=C_{0}+C_{1}$ reveals that distributional shift affects the two classes asymmetrically ($2.03\times$ for survival vs. $1.30\times$ for mortality): the larger OoD signal resides in the non-critical class, so $C_{1}$-only yields AUROC $0.740$, well below $\sum_{k}C_{k}$ at $0.815$. On FashionMNIST the pattern is the opposite: all ten per-class $C_{k}$ values increase uniformly from ID to OoD (Appendix E.4, Figure 11), consistent with KMNIST inducing a structural rather than semantic shift that does not concentrate on any subset of classes. Together, these two cases illustrate a task-dependent trade-off: critical-class targeting improves selective prediction (Section 3) but all-class aggregation is necessary for OoD detection, where the locus of distributional shift is not known a priori and can be either asymmetric (MIMIC-III) or uniform (FashionMNIST). Full per-class results in Appendix E.3.
@@ 5.1 Setup
We use Fashion-MNIST (K=10K{=}10, 60,000 images) and CIFAR-10 (K=10K{=}10, 50,000 images). For each dataset we inject symmetric label noise at rates $\alpha\in\{0.1,0.2,0.3,0.4,0.5\}$: a fraction $\alpha$ of training labels is replaced uniformly at random, increasing aleatoric uncertainty without altering the input distribution.
@@
Both tasks use fully Bayesian networks trained end-to-end from scratch with low-rank Gaussian posteriors. Fashion-MNIST: two hidden layers ($400$ units, rank $15$), S=50S{=}50 MC samples. CIFAR-10: Conv $32{\to}64{\to}128$ with global average pooling and a low-rank Bayesian head (rank $32$), S=50S{=}50 MC samples. Full architectural details are in Appendix F.1.
@@ 5.2 Results
The transfer-learning column in Table 4 already isolates training regime as a causal factor; within a single regime the posterior family matters too. On Fashion-MNIST, where all three model types are trained end-to-end, the low-rank model achieves a mean $|R_{\mathrm{rel}}|$ below $0.01$ for both metrics, while the full-rank model reaches $0.11$ (MI) / $0.10$ ($\sum_{k}C_{k}$) and MC dropout $0.10$ / $0.06$ respectively (Appendix F.2). Under transfer learning the ordering is preserved across both CIFAR-10 and CIFAR-100: low-rank << full-rank << MC dropout, with MC dropout ratios exceeding $0.58$ on CIFAR-100 at $\alpha{=}0.1$. This suggests that the high entanglement reported by mucsanyi2024benchmarking reflects not only limitations of the decomposition formula but also the quality of the underlying posterior approximation: models that do not propagate uncertainty through the full network produce predictive distributions whose variance structure is poorly suited to any epistemic measure, whether MI or $\sum_{k}C_{k}$. Across all end-to-end configurations (Figure 15), $\sum_{k}C_{k}$ achieves lower $|R_{\mathrm{rel}}|$ than MI in 19 of 20 conditions (4 models $\times$ 5 noise levels). Under transfer learning the advantage is partial, with MI prevailing where baseline inflation $\sum_{k}C_{k}\,/\,\text{MI}$ exceeds $1.6\times$ (Table 22), i.e. precisely where the Taylor approximation is least accurate. Full transfer-learning results and an analysis of the rank constraint’s effect on uncertainty scale are in Appendix F.2.
@@
On CIFAR-100, the $1/\mu_{k}$ normalisation amplifies contributions from low-probability classes, inflating $\sum_{k}C_{k}$ relative to MI by $1.17\times$ for the low-rank model and $1.89\times$ for MC dropout. The $\mathcal{O}(K^{2})$ scaling analysis and mitigation strategies are detailed in Appendix F.3.
@@ 6 Discussion and Conclusion
We presented a per-class epistemic uncertainty decomposition $\boldsymbol{C}(x)$ derived from a second-order Taylor expansion of MI, where the $1/\mu_{k}$ normalisation arises naturally from the entropy Hessian and counteracts boundary suppression. The decomposition reduces to a simple ratio of variance to mean per class, adds negligible cost to any posterior sampling pipeline, and recovers scalar MI by summation.
@@
Several limitations exist. The additive link $\sum_{k}C_{k}\approx\mathcal{I}$ loosens under high skewness over low-probability classes, since the third-order remainder scales as $1/\mu_{k}^{2}$; $\rho_{k}$ diagnoses but does not correct this, and the CBEC fallback requires domain knowledge to define class partitions. The $1/\mu_{k}$ normalisation introduces $\mathcal{O}(K^{2})$ scaling of the aggregate, necessitating truncation or reweighting in high-cardinality settings, though the per-class vector itself remains fully interpretable. All models use MFVI, low-rank Gaussian VI, MC dropout or deep ensembles. Characterising how $C_{k}$ behaves under other approximate inference schemes such as Laplace approximations [daxberger2021laplace] or stochastic weight averaging [maddox2019simple] is a natural next step.
The decomposition is not tied to any specific posterior family or architecture. Future work includes extending $\boldsymbol{C}(x)$ to structured prediction tasks where class identity is less cleanly defined, integrating it into active learning acquisition functions that target specific classes, and combining low-rank ensembles with per-class attribution for richer uncertainty profiles. More fundamentally, making end-to-end Bayesian training more practical strikes us as a more promising path than refining post-hoc corrections, and the per-class decomposition is designed to complement that direction. [^4] Code and trained models are available in the supplementary material.

### v2

@@ 2 Per-Class Epistemic Uncertainty
We decompose MI into a $K$-dimensional vector $\boldsymbol{C}(x)$ via second-order Taylor expansion of the entropy. Each component $C_{k}$ quantifies class $k$’s epistemic contribution, and the sum approximates MI.
@@
with diagonal elements $\mathrm{Var}[p_{k}]=\frac{1}{S-1}\sum_{s}(p_{k}^{(s)}-\mu_{k})^{2}$. Throughout, we rely on the standard decomposition
@@ 2.1 From Scalar MI to a Per-Class Decomposition
The Hessian is diagonal and negative semidefinite on the interior of the simplex. The diagonal Hessian is what makes a per-class decomposition possible: there are no cross-class terms at second order, so the quadratic approximation splits cleanly over classes.
Theorem 2.2 (MI Approximation 222smith2018understanding also exploit a Taylor expansion to relate softmax variance to MI, but expand the logarithm $\log p_{k}$ rather than the entropy function $H(\boldsymbol{p})$ itself (their Section 3.2, Eq. 10). This yields a class-averaged variance $\frac{1}{K}\sum_{k}\mathrm{Var}[p_{k}]$ with equal weights per class, whereas expanding the entropy via its Hessian yields the curvature-weighted sum $\frac{1}{2}\sum_{k}\mathrm{Var}[p_{k}]/\mu_{k}$, in which the $1/\mu_{k}$ factor enables per-class attribution.).
@@
The simplex constraint forces $\mathrm{Cov}[\boldsymbol{p}]$ to have rank at most K1K{-}1, coupling the $K$ components; this does not invalidate the trace formula. See Appendix B.1 for proof.
@@ 2.3 Why Variance Alone Fails: Boundary Suppression
which approaches $\frac{1}{2}$ (not zero) as $\mu_{k}\to 0$. While $\mathrm{Var}[p_{k}]$ is crushed to zero near the simplex boundary, $C_{k}$ retains a non-vanishing upper bound. See Appendix B.2.
@@ 3.1 Experimental Setup
Uncertainty is estimated from $S=30$ stochastic forward passes through the posterior. Our primary safety metric is the critical false negative rate (Critical FNR, Appendix C.8): among critical-class samples retained at a given coverage level, the fraction misclassified as safe. We summarise performance across coverage levels via the area under the selective risk curve (AUSC) [elyaniw2010], with 95% CIs from 200 bootstrap resamples. On the test set, the Bayesian model achieves accuracy $0.8$ and quadratic weighted kappa $0.65$; per-class performance in Appendix D.3.
@@ 3.3 Results
Table 2 reports AUSC and Critical FNR at 80% coverage with bootstrap CIs. $C_{\text{crit\_max}}$ achieves the lowest AUSC ($0.285$), reducing the area under the selective risk curve by 34.7% relative to MI ($0.436$, $p<0.005$) and by 56.2% relative to Sale_EU_crit ($0.650$, $p<0.005$). At the clinically relevant 80% coverage operating point, $C_{\text{crit\_max}}$ achieves FNR $0.302$, compared to $0.339$ for MI and $0.409$ for Sale_EU_crit. Figure 1 visualises these results: $C_{\text{crit\_max}}$ dominates across the full coverage range (left), with non-overlapping bootstrap interquartile ranges against scalar baselines (right). The improvement over raw variance is especially striking. Sale_EU_crit and Var_crit (AUSC $0.650$ and $0.606$) perform worse than entropy ($0.604$), confirming that boundary suppression (Lemma 2.4) severely confounds unnormalised variance for critical classes with $\mu_{k}\approx 0.06$. The $1/\mu_{k}$ normalisation in $C_{k}$ corrects precisely this: Grade 3 has mean probability $\sim$0.06, giving entropy curvature $\sim$17, so even small variance there carries large information-theoretic weight. This normalisation benefit is robust to the choice of inference method: a deep ensemble [lakshminarayanan2017simple] of five members yields $C_{\text{crit\_max}}$ AUSC $0.390$ vs. $0.447$ for Sale_EU_crit ($12.9\%$ reduction; Appendix D.12).
@@
Robustness across inference methods.
To test robustness to the posterior approximation, we repeat the evaluation using standard MC dropout [gal2016dropout] (p=0.3p{=}0.3, S=30S{=}30). The ranking changes markedly: CBEC achieves the lowest AUSC ($0.197\pm 0.012$, winning 100% of bootstrap samples), a 53.6% reduction over MI ($0.425\pm 0.024$), while $C_{\text{crit\_max}}$ drops to $0.419$. MC dropout is known to produce regions of spurious confidence where the approximate posterior underestimates uncertainty [smith2018understanding]. In our framework, this inflates $\rho_{k}$ for critical classes, degrading the Taylor approximation underlying $C_{k}$ while leaving CBEC’s correlation gate intact. A deep ensemble of five members confirms the pattern: CBEC again achieves the lowest AUSC ($0.223$), even though $C_{k}$ reliability remains high ($\rho_{k}<0.3$ for $85\%$ of Grade 3 samples), indicating that the correlation gate captures cross-boundary structure beyond what single-class targeting provides. This confirms the intended complementarity: $C_{\text{crit\_max}}$ is the primary metric under well-calibrated posteriors; CBEC provides a robust alternative across inference regimes. Full results in Appendix D.11 and D.12.
@@ 4.1 Setup
[…]
@@ 4.2 Results
Table 3 summarises OoD detection performance. $\sum_{k}C_{k}$ achieves the highest AUROC on both datasets: $0.735$ on FashionMNIST versus $0.724$ for MI, and $0.815$ on MIMIC-III versus $0.802$ for MI. The OoD-to-ID ratio column reveals the mechanism: on FashionMNIST, $\sum_{k}C_{k}$ amplifies the distributional shift signal to $6.43\times$ versus $5.92\times$ for MI. On FashionMNIST, this effect is predicted by Lemma 2.9: the gap $\sum_{k}C_{k}-\text{MI}\approx\frac{1}{6}\sum_{k}m_{3,k}/\mu_{k}^{2}$ grows with posterior skewness, and OoD samples produce higher skewness diagnostic $\rho_{k}$ than ID samples across all ten classes (Appendix E.6, Figure 14). On MIMIC-III the picture is more nuanced: the survival class (k=0k{=}0) follows the expected pattern ($\rho_{0}^{\text{OOD}}>\rho_{0}^{\text{ID}}$, median shift $43\%$), but the mortality class (k=1k{=}1) reverses because near-zero $\mu_{1}$ for newborns collapses $\mathrm{Var}[p_{1}]$ faster than $1/\mu_{1}$ can compensate, concentrating the third-order signal entirely in class 0 (Appendix E.6).
The unnormalised baseline $EU_{\text{var}}$ underperforms MI on both datasets despite matching its ratio on FashionMNIST ($5.92\times$): without mean-normalisation, raw variance $\sigma^{2}_{k}$ is bounded above by $\mu_{k}(1-\mu_{k})$, compressing the dynamic range of uncertainty scores. On MIMIC-III $EU_{\text{var}}$ achieves the highest ratio ($1.71\times$) yet the lowest AUROC among epistemic measures. A high mean ratio does not imply good separability: AUROC depends on the full distributional overlap, and the un-normalised variance sum concentrates both ID and OoD scores near zero with a shared heavy tail, so the numerically large mean shift is swamped by within-group spread (Appendix E.5, Figure 12). This pattern is consistent across inference methods: a deep ensemble replicates the same ranking ($\sum_{k}C_{k}{=}0.753>\text{MI}{=}0.752>EU_{\text{var}}{=}0.750$) on MIMIC-III, with $EU_{\text{var}}$ again achieving the highest ratio ($2.72\times$) and the lowest AUROC (Appendix E.7).
@@
On the binary MIMIC-III task, the decomposition $\sum_{k}C_{k}=C_{0}+C_{1}$ reveals that distributional shift affects the two classes asymmetrically ($2.15\times$ for survival vs. $1.30\times$ for mortality): the larger OoD signal resides in the non-critical class, so $C_{1}$-only yields AUROC $0.740$, well below $\sum_{k}C_{k}$ at $0.815$. On FashionMNIST the pattern is the opposite: all ten per-class $C_{k}$ values increase uniformly from ID to OoD (Appendix E.4, Figure 11), consistent with KMNIST inducing a structural rather than semantic shift that does not concentrate on any subset of classes. Together, these two cases illustrate a task-dependent trade-off: critical-class targeting improves selective prediction (Section 3) but all-class aggregation is necessary for OoD detection, where the locus of distributional shift is not known a priori and can be either asymmetric (MIMIC-III) or uniform (FashionMNIST). Full per-class results in Appendix E.3.
@@ 5.1 Setup
We use Fashion-MNIST (K=10K{=}10, 60,000 images) and CIFAR-10 (K=10K{=}10, 50,000 images) [krizhevsky2009learning]. For each dataset we inject symmetric label noise at rates $\alpha\in\{0.1,0.2,0.3,0.4,0.5\}$: a fraction $\alpha$ of training labels is replaced uniformly at random, increasing aleatoric uncertainty without altering the input distribution.
@@
We compare end-to-end Bayesian training with transfer learning using low-rank Gaussian posteriors (S=50S{=}50 MC samples). Fashion-MNIST: fully Bayesian; two hidden layers ($400$ units, rank $15$). CIFAR-10: (i) end-to-end Conv $32{\to}64{\to}128$, rank $32$; (ii) frozen ResNet-18 with Bayesian head. Full details, alternative posteriors, and CIFAR-100 results in Appendix F.1.
@@ 5.2 Results
Table 4 identifies training regime as a key factor. Under end-to-end training, posterior expressiveness tracks entanglement: on Fashion-MNIST the low-rank model achieves a mean $|R_{\mathrm{rel}}|$ below $0.01$ for both metrics, while the full-rank model reaches $0.11$ (MI) / $0.10$ ($\sum_{k}C_{k}$) and MC dropout $0.10$ / $0.06$ respectively (Appendix F.2). Transfer learning disrupts this hierarchy: even the low-rank model produces $|R_{\mathrm{rel}}|$ exceeding $0.73$ on CIFAR-10, an order of magnitude worse than any end-to-end configuration, because the frozen backbone supplies features never optimised for calibrated predictive variance. This suggests that the high entanglement reported by mucsanyi2024benchmarking reflects the training regime as much as the decomposition formula. When uncertainty is not propagated through the full network, none of the three posterior families we evaluate can compensate, motivating the skewness diagnostic of Section 2.5 to flag such configurations. Across end-to-end settings (Figure 15), $\sum_{k}C_{k}$ achieves lower $|R_{\mathrm{rel}}|$ than MI in 19 of 20 conditions. Under transfer learning both metrics degrade and MI edges ahead, coinciding with baseline ratios $\sum_{k}C_{k}\,/\,\text{MI}$ above roughly $1.6\times$ (Table 22), i.e. where the Taylor residual is largest. Full transfer-learning results and an analysis of the rank constraint’s effect on uncertainty scale appear in Appendix F.2.
@@
On CIFAR-100 [krizhevsky2009learning], the $1/\mu_{k}$ normalisation amplifies contributions from low-probability classes, inflating $\sum_{k}C_{k}$ relative to MI by $1.17\times$ for the low-rank model and $1.89\times$ for MC dropout. The $\mathcal{O}(K^{2})$ scaling analysis and mitigation strategies are detailed in Appendix F.3.
@@ 6 Discussion and Conclusion
We presented a per-class epistemic uncertainty decomposition $\boldsymbol{C}(x)$ derived from a second-order Taylor expansion of MI, where the $1/\mu_{k}$ normalisation arises naturally from the entropy Hessian and counteracts boundary suppression. The decomposition reduces to a simple ratio of variance to mean per class, adds negligible cost to any posterior sampling pipeline, and closely tracks scalar MI by summation.
@@
We note several limitations. The additive link $\sum_{k}C_{k}\approx\mathcal{I}$ loosens under high skewness over low-probability classes, since the third-order remainder scales as $1/\mu_{k}^{2}$; $\rho_{k}$ diagnoses but does not correct this, and the CBEC fallback requires domain knowledge to define class partitions. The $1/\mu_{k}$ normalisation introduces $\mathcal{O}(K^{2})$ scaling of the aggregate, necessitating truncation or reweighting in high-cardinality settings, though the per-class vector itself remains fully interpretable. All models use MFVI, low-rank Gaussian VI, MC dropout, or deep ensembles. Characterising how $C_{k}$ behaves under other approximate inference schemes such as Laplace approximations [daxberger2021laplace] or stochastic weight averaging [maddox2019simple] is a natural next step.
The decomposition is not tied to any specific posterior family or architecture. Future work includes extending $\boldsymbol{C}(x)$ to structured prediction tasks where class identity is less cleanly defined, integrating it into active learning acquisition functions that target specific classes, and combining low-rank ensembles with per-class attribution for richer uncertainty profiles. More fundamentally, making end-to-end Bayesian training more practical and scalable strikes us as a more promising path than refining post-hoc corrections, and the per-class decomposition is designed to complement that direction. [^4] Code and trained models are available in the supplementary material.